Two digits are selected at random from the digits 1 through 9





1
Expert's answer
2021-12-20T16:14:29-0500

If the sum is even, find the probability that both the numbers are odd.


Solution:


Let A=Getting two odd numbers 

B=Getting the sum as an even number.

Required probability:


"P(A|B)=\frac{P(A\cap B)}{P(B)}=\frac{C^2_5/C^2_9}{(C^2_4+C^2_4)/C^2_9}=\frac{C^2_5}{C^2_4+C^2_5}=10/16=5/8"